what is 41% of 300 students.

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

what is 41% of 300 students?
Ans: It is 41 out of each 100 students.
41%*300 = (41/100)300 = 41(300/100) = 41*3 = 123 students
----------------
Cheers,
Stan H.
-------------------

Answer by rfer(16322)

300*0.41=123